# ICNN-enhanced 2SP: Leveraging input convex neural networks for solving two-stage stochastic programming

### v1

@@ 3 Convexity analysis
In summary, for a broad class of 2SP problems, the recourse function $\mathcal{Q}(x)$ inherits a piecewise convex structure, even accommodating relatively challenging instances with strategic integer decisions in the first stage and operational adjustments (continuous) or logical switches (binary) in recourse actions. Such structural regularity makes approximation via ICNNs advantageous, as they inherently preserve convexity while learning surrogate models of $\mathcal{Q}(x)$, enabling the embedding of the ICNN surrogates into the original 2SP for efficient solving.
@@ 4 Methodology
This section presents our ICNN-enhanced framework for 2SP that directly leverages the convexity properties of $\mathcal{Q}(x)$. By exploiting this inherent convexity, we employ ICNNs as recourse function approximators and embed this surrogate into the first-stage problem to obtain a reformulated 2SP without introducing integer variables.
@@ 4.1 Base architecture
Building on the extensive form (2.1), Neur2SP constructs a surrogate model $\hat{\mathcal{Q}}(x)$ to explicitly model the expected second-stage cost $\mathcal{Q}(x)$ through two sequential transformations:
(i) Scenario Encoding: consists of the $x$-independent processing of each scenario $\xi_{s}$ through a shared network $\Psi^{1}$ to transform each scenario into a latent feature vector, followed by mean-aggregation and final encoding via another network $\Psi^{2}$:
@@
(ii) Decision Mapping: represents a conventional ReLU NN $\Phi$ that combines first-stage decisions $x$ with scenario encoding $\xi_{\lambda}$ to predict $\mathbb{E}_{\xi}[Q(x,\xi)]$:
@@
While effective for general 2SP, this architecture becomes suboptimal for convex problems due to its MIP-based surrogate embeddings. The embedding of ReLU NNs as MIP constraints within optimisation problems is detailed in Appendix A. In our approach, we retain Neur2SP’s core architecture for scenario processing while being specialised for convex 2SP models. The key difference lies in $\Phi$’s architecture: $\Phi$ is formulated as an ICNN since $Q(x,\xi_{s})$ is convex in $x$ (per Section 3).
@@ 4.2 Surrogate modelling
$\min_{\Theta}\sum_{i}\Biggl{(}\Phi_{\text{ICNN}}\biggl{(}x^{(i)},\Psi^{2}\Bigl{(}\frac{1}{|S|}\sum_{s=1}^{|S|}\Psi^{1}(\xi_{s}^{(i)})\Bigl{)}\biggr{)}-\mathbb{E}_{\xi}[Q(x^{(i)},\xi)]\Biggr{)}^{2},$ (11)
where $\Theta=\{W_{k},S_{k},b_{k}\}\cup\{\Psi^{1},\Psi^{2}\}$ represents the complete set of trainable parameters. Importantly, once trained, the model generalises to arbitrary second-stage scenario sets without requiring retraining, eliminating scenario set-specific training overhead.
@@ 5.1 Experimental setup
All experiments were conducted on a Triton computing cluster maintained by the Aalto Science-IT project, utilising an Intel Xeon Gold 6148 CPU and an NVIDIA Tesla H100 GPU with 64 GB of RAM. Our code extends Neur2SP’s codebase (Patel and Dumouchelle, 2022) by augmenting it to include all necessary ICNN-related functionalities. The implementation uses Python 3.11.5 with Gurobi 11.0.2 (Gurobi Optimization, LLC, 2024) as the optimisation solver, PyTorch 2.0.1 (Paszke et al., 2019) and scikit-learn 1.3.0 (Pedregosa et al., 2011) for NN training. Full reproducibility materials are provided at  https://github.com/Lycle/ICNN2SP.
We evaluate our method on three canonical 2SP problems which are also taken as benchmark problems in Neur2SP (Dumouchelle et al., 2022): the capacitated facility location problem (CFLP) (Cornuejols et al., 1991), the stochastic server location problem (SSLP) (Ntaimo and Sen, 2005), and the investment problem (INVP) (Schultz et al., 1998). Problem instances are denoted by type followed by structural parameters, with the final parameter indicating scenario count. Detailed descriptions are provided in Appendix B. The EF serves as the ground-truth baseline (considering a 3-hour time limit). We compare both ICNN-enhanced 2SP and Neur2SP against EF to contextualise performance.
To ensure direct comparability with Neur2SP, we replicate their instance generation protocol, including uncertainty distributions and scenario sampling methods. We generate 5,000 first-stage solutions by uniformly sampling the feasible region. For each solution, the expected recourse cost is approximated by solving its second-stage problem for a random subset of scenarios, with the number of scenarios per solution randomly selected up to 100, which balances computational efficiency with the diversity of scenarios. Further details on data generation are provided in Appendix C. For both methods, we conduct a random search over 100 hyperparameter configurations for NN architecture selection, detailed in Appendix D. To mitigate solver-induced variability, Gurobi’s parameters are fixed across all runs, as detailed in Appendix E.
@@ 5.2 Results and discussion
Tables 1 to 4 report our experimental results. In addition, we include supplementary results with the objective values in Appendix F and results for the SSLP instances from the Stochastic Integer Programming Test Problem Library (SIPLIB) (Ahmed et al., 2015) in Appendix G. We also analyse the variable counts in Appendix H to compare formulation size.
Both methods share the same structural design for Scenario Encoding ($\Psi^{1}$ and $\Psi^{2}$) while differing in the architectural choice of Decision Mapping ($\Phi$): our method uses ICNNs while Neur2SP employs conventional ReLU NNs, which constitutes the primary source of training performance differences. As shown in Table 1, both methods exhibit training times within the same order of magnitude, with ICNNs requiring only marginally longer training due to their convexity constraints. This modest overhead is counterbalanced by their comparable or superior validation accuracy, achieving lower mean absolute error (MAE) in 7 out of 10 problem instances. These results demonstrate that the convexity constraints do not significantly compromise the expressive power of the NN while presenting mathematical properties that enable more efficient downstream optimisation.
@@
Tables 24 reveal consistent performance advantages for ICNN-enhanced 2SP across problem classes. For CFLP, our method demonstrates superior scalability, achieving 1.3$\times$ speedups for small instances and up to 100$\times$ faster solving times for more complex problem class (CFLP_50_50) when compared to Neur2SP, while maintaining comparable or better solution quality. For SSLP, our approach achieves 2$\times$ speedups with identical optimality gaps for SSLP_5_25, a substantial gap reduction with 1.1-1.2$\times$ speedups for SSLP_15_45, and 3$\times$ faster solution times with matching solution quality for SSLP_10_50 instances. For INVP, while absolute solving time differences are marginal ($\leq$0.03 s), ICNN-enhanced 2SP consistently delivers superior solution quality across most instances, often with considerable optimality gap improvements over Neur2SP.
Table 2: CFLP results (averaged over 11 instances per row; EF averages include only instances where EF found feasible solutions within 3h). Gap: relative objective difference (%) of ICNN-enhanced 2SP (ICNN-EF) and Neur2SP (NN-EF) compared to EF; negative values indicate better solutions for minimisation. Solving time: runtime (s); $\sim$10,800 indicates time limit reached. Bold indicates better performance.
Gap (%) Solving time (s)
Problem ICNN-EF NN-EF ICNN NN EF
CFLP_10_10_100 0.57 0.51 0.29 0.39 8,491.85
CFLP_10_10_500 0.15 0.00 0.29 0.39 10,800.05
CFLP_10_10_1000 -0.02 -0.27 0.29 0.39 10,800.29
CFLP_25_25_100 2.19 3.37 0.39 10.31 10,800.05
CFLP_25_25_500 1.74 2.63 0.40 10.37 10,800.12
CFLP_25_25_1000 0.51 1.48 0.40 10.07 10,800.33
CFLP_50_50_100 3.36 18.17 0.81 84.07 10,800.63
CFLP_50_50_500 -9.82 3.06 0.84 84.32 10,800.16
CFLP_50_50_1000 -12.82 -0.28 0.84 84.39 10,800.29
Table 3: SSLP results (averaged over 12 instances per row). All settings and metric definitions follow Table 2.
Gap(%) Solving time (s)
Problem ICNN-EF NN-EF ICNN NN EF
SSLP_5_25_50 0.12 0.12 0.13 0.25 1.58
SSLP_5_25_100 0.02 0.02 0.12 0.25 5.18
SSLP_15_45_5 2.65 18.09 0.16 0.20 1.94
SSLP_15_45_10 2.57 18.38 0.18 0.20 32.74
SSLP_15_45_15 2.11 18.72 0.20 0.21 605.22
SSLP_10_50_50 0.00 0.00 0.27 0.76 9,007.88
SSLP_10_50_100 0.00 0.00 0.26 0.76 10,800.04
SSLP_10_50_500 0.00 0.00 0.28 0.77 10,800.21
SSLP_10_50_1000 -2.22 -2.22 0.28 0.77 10,803.79
SSLP_10_50_2000 -102.68 -102.68 0.29 0.78 10,800.14
@@
Gap (%) Solving time (s)
Problem ICNN-EF NN-EF ICNN NN EF
INVP_B_E_4 11.06 10.80 0.06 0.15 0.00
INVP_B_E_9 5.60 4.31 0.07 0.10 0.01
INVP_B_E_36 3.28 4.58 0.06 0.10 0.05
INVP_B_E_121 3.07 2.85 0.06 0.09 1.32
INVP_B_E_441 3.10 3.18 0.06 0.10 125.60
INVP_B_E_1681 0.98 0.21 0.07 0.10 10,800.00
INVP_B_E_10000 -1.59 -1.64 0.07 0.11 10,808.33
INVP_B_H_4 8.81 8.81 0.06 0.08 0.01
INVP_B_H_9 5.04 5.04 0.07 0.09 0.01
INVP_B_H_36 1.61 1.61 0.07 0.08 1.18
INVP_B_H_121 1.77 1.77 0.06 0.08 60.09
INVP_B_H_441 2.13 2.13 0.06 0.09 322.05
INVP_B_H_1681 0.69 0.69 0.06 0.09 10,800.01
INVP_B_H_10000 -2.72 -2.72 0.07 0.10 10,808.33
INVP_I_E_4 4.78 17.32 0.05 0.08 0.00
INVP_I_E_9 1.07 7.01 0.04 0.07 0.01
INVP_I_E_36 0.37 3.98 0.06 0.07 0.03
INVP_I_E_121 0.17 2.88 0.06 0.08 0.39
INVP_I_E_441 0.11 3.22 0.04 0.07 39.73
INVP_I_E_1681 0.42 3.39 0.05 0.07 10,800.00
INVP_I_E_10000 0.22 3.50 0.05 0.08 10,800.09
INVP_I_H_4 13.78 13.78 0.07 0.06 0.01
INVP_I_H_9 5.85 9.12 0.05 0.07 0.01
INVP_I_H_36 3.69 4.97 0.06 0.07 0.40
INVP_I_H_121 3.75 4.01 0.05 0.07 7.27
INVP_I_H_441 1.27 3.15 0.05 0.07 10,800.00
INVP_I_H_1681 -0.41 -0.34 0.05 0.07 10,800.01
INVP_I_H_10000 0.27 0.25 0.06 0.08 10,800.03
Overall, ICNN-enhanced 2SP demonstrates computational advantages across various problem settings, with solution quality matching or exceeding Neur2SP in 80% of cases. While training requires modest overhead (10% longer on average), this upfront investment enables acceleration during optimisation through LP-driven formulation.
6 Conclusions
This work establishes ICNNs as an alternative framework for solving convex 2SP, offering an effective pathway to reconcile computational efficiency with solution quality. Our framework, ICNN-enhanced 2SP, integrates ICNNs as surrogates for the expected recourse function, showing that its architectural convexity constraints can be strategically harnessed rather than being limitations. Across diverse problem classes, ICNN-enhanced 2SP outperforms existing MIP-based methodologies, with its computational advantages becoming more pronounced as problem scale increases. The computational efficiency makes our approach viable for time-critical decision-making applications requiring frequent resolving, such as real-time power dispatch and production scheduling, where solution time directly impacts operational feasibility, and MIP-based alternatives would be prohibitively expensive.
Acknowledgments and Disclosure of Funding
Appendix A MIP reformulations
Consider a conventional ReLU NN with $L+1$ layers (numbered from 00 to $L$) to build the surrogate model of an underlying function $f$, which can be approximated and effectively replaced using this ReLU NN surrogate $\hat{f}$, constructed based on dataset $\mathcal{D}$. For the input layer, we have $y^{0}=x$. For the output layer we have $y^{L}=\hat{f}_{\mathcal{D}}(x)$. The activation of neurons $i=1,\dots,N_{l}$ in the hidden layer $l=1,\dots,L-1$ are calculated as
$L_{i}^{l}\leq w_{i}^{l^{\intercal}}y_{l-1}+b^{l}_{i}\leq U_{i}^{l}.$ (13)
Following Fischetti and Jo (2018), the ReLU operator can be encoded into mixed-integer linear constraints by introducing slack variables $s^{l}_{i}$ and binary variables $z^{l}_{i}$ for $i=1,\dots,N_{l},\ l=1,\dots,L-1$, which is expressed as:
$y_{i}^{l}-s_{i}^{l}=w_{i}^{l^{\intercal}}y^{l-1}+b_{i}^{l},$ (14)
$0\leq y_{i}^{l}\leq U_{i}^{l}z_{i}^{l},$ (15)
$0\leq s_{i}^{l}\leq-L_{i}^{l}\left(1-z_{i}^{l}\right),$ (16)
$z_{i}^{l}\in\left\{0,1\right\}.$ (17)
Combining these constraints with input and output layer bounds
$L^{0}\leq y^{0}\leq U^{0},$ (18)
$L^{L}\leq y^{L}=w^{L^{\intercal}}y^{L-1}+b^{L}\leq U^{L},$ (19)
Appendix B Stochastic programming problems
This appendix provides brief descriptions of the three benchmark problems employed in our experimental evaluation. These are the same problems utilised in Dumouchelle et al. (2022), facilitating direct comparison between our ICNN-enhanced approach and the Neur2SP methodology.
B.1 Capacitated facility location problem (CFLP)
The CFLP addresses the challenge of selecting optimal facility locations to service customer demand at minimal cost. In its stochastic formulation, facility decisions precede knowledge of actual demand patterns. Our implementation adapts the case from Cornuejols et al. (1991) to the stochastic setting by preserving the original cost structures and capacity parameters while stochastically sampling demand values. To guarantee relatively complete recourse, we incorporate penalty variables for potential demand shortfalls. Throughout our experimental analysis, we denote a CFLP instance with $n$ facilities, $m$ customers, and $s$ scenarios as CFLP_$n$_$m$_$s$.
B.2 Stochastic server location problem (SSLP)
The SSLP focuses on optimal server deployment and client assignment under uncertain service requests. Binary first-stage variables determine server placement, while second-stage variables manage client-server assignments after observing which clients require service. Our experimental protocol utilises the standard benchmark instances from SIPLIB Ahmed et al. (2015) as documented in Ntaimo and Sen (2005). For consistent reference, we denote an SSLP instance with $n$ potential server locations, $m$ possible clients, and $s$ scenarios as SSLP_$n$_$m$_$s$.
B.3 Investment problem (INVP)
The INVP consists of continuous first-stage investment decisions followed by discrete second-stage profit-generating activities. We implement Example 7.3 from Schultz et al. (1998), with two first-stage variables bounded by $[0,5]$ and four binary second-stage variables. Scenarios derive from two random variables uniformly distributed over $[5,15]$. For consistency, we reformulate it as a minimisation problem. In our experiments, we denote instances as INVP_$v$_$t$_$s$, where $v$ indicates second-stage variable type (B for binary, I for integer), $t$ indicates technology matrix type (E for identity, H for $[[2/3,1/3],[1/3,2/3]]$), and $s$ indicates the number of scenarios.
Appendix C Data generation
(IC)NN
Problem Time per sample (s) Total time (s)
CFLP_10_10 0.32 1,620.07
CFLP_25_25 2.22 11,096.88
CFLP_50_50 2.19 10,945.11
SSLP_5_25 0.84 4,182.60
SSLP_15_45 2.49 12,438.02
SSLP_10_50 2.33 11,649.44
INVP_B_E 1.77 8,855.89
INVP_B_H 1.43 7,157.07
INVP_I_E 1.48 7,420.66
INVP_I_H 1.04 5,194.90
Appendix D Model selection
For model selection, we conducted a random search over 100 configurations for each problem instance following Dumouchelle et al. (2022). Both the ICNN-enhanced 2SP and Neur2SP methods were subjected to identical hyperparameter sampling from the distribution space outlined in Table 6, with training proceeding for 2,000 epochs across all configurations.
Parameter Sampling range
Batch size $\{16,32,64,128\}$
Learning rate $[1e^{-5},1e^{-1}]$
L1 weight penalty $[1e^{-5},1e^{-1}]$
L2 weight penalty $[1e^{-5},1e^{-1}]$
Optimiser {Adam, Adagrad, RMSprop}
Dropout $[0,0.5]$
ReLU hidden dimension $\{64,128,256,512\}$
Embed hidden dimension 1 $\{64,128,256,512\}$
Embed hidden dimension 2 $\{16,32,64,128\}$
Embed hidden dimension 3 $\{8,16,32,64\}$
Appendix E Solver parameter settings
Gurobi’s automatic parameter selection introduces variability in results due to divergent solution strategies rather than formulation differences. To isolate methodological comparisons, we fix key Gurobi parameters (Table 7), standardising the solving environment across formulations. See Gurobi Optimization, LLC (2025) for details.
Table 7: Solver parameter settings in the experiments.
Parameter Value Description
TimeLimit 10800 Limit total runtime (3 h)
Threads 1 Thread count
Presolve 0 Disable presolve
Heuristics 0 Minimise time spent in MIP heuristics
NodeMethod 1 Use dual simplex for MIP node relaxations
Cuts 1 Moderate cut generation
MIPFocus 2 Focus on proving optimality
Appendix F Objective results
This section presents a detailed comparison of objective function values obtained through different solution approaches. We report the objective values of the EF formulation as the baseline, which represents the best solutions found within the time limit. Additionally, we report two sets of objectives for solutions obtained via surrogate embedding reformulations: the “true objective” and the “approximate objective”. Tables 8 through 10 present these comparative values. The results demonstrate that the ICNN-enhanced approach consistently achieves solution quality on par with or better than conventional NN approaches while maintaining computational advantages discussed in the main paper.

### v2

@@ 3.4 Practical implications for ICNN surrogates
In summary, for a broad class of 2SP problems, the recourse function $\mathcal{Q}(x)$ inherits a piecewise convex structure, even accommodating relatively challenging instances with strategic integer decisions in the first stage and operational adjustments (continuous) or logical switches (binary) in recourse actions. Such structural regularity makes approximation via ICNNs advantageous, as they inherently preserve convexity while learning surrogate models of $\mathcal{Q}(x)$, enabling the embedding of the ICNN surrogates into the original 2SP for efficient solving. In cases where convexity cannot be analytically guaranteed, a practical approach is first to test whether the recourse function exhibits approximate convexity, for example, by training a lightweight ICNN and evaluating whether its predictive accuracy is acceptable. A symmetric check for approximate concavity is also possible by negating the training labels and assessing whether an ICNN fits well under this transformation.
@@ 4 Methodology
This section presents our ICNN-enhanced framework for 2SP that directly leverages the convexity properties of $\mathcal{Q}(x)$. The ICNN surrogate of the recourse function is embedded into the first-stage problem to obtain a reformulated 2SP without introducing integer variables.
@@ 4.1 Base architecture
Building on the extensive form (3), Neur2SP constructs a surrogate model $\hat{\mathcal{Q}}(x)$ to explicitly model the expected second-stage cost $\mathcal{Q}(x)$ through two sequential transformations:
  1. (i)
    Scenario Encoding: consists of the $x$-independent processing of each scenario $\xi_{s}$ through a shared network $\Psi^{1}$ to transform each scenario into a latent feature vector, followed by mean-aggregation and final encoding via another network $\Psi^{2}$:
@@
  2. (ii)
    Decision Mapping: represents a conventional ReLU NN $\Phi$ that combines first-stage decisions $x$ with scenario encoding $\xi_{\lambda}$ to predict $\mathbb{E}_{\xi}[Q(x,\xi)]$:
@@
While effective for general 2SP, this architecture becomes suboptimal for convex problems due to its MIP-based surrogate embeddings. The embedding of ReLU NNs as MIP constraints within optimisation problems is detailed in B. In our approach, we retain Neur2SP’s core architecture for scenario processing while being specialised for convex 2SP models. The key difference lies in $\Phi$’s architecture: $\Phi$ is formulated as an ICNN when $Q(x,\xi_{s})$ is convex in $x$ (per Section 3).
@@ 4.2 Surrogate modelling
$\min_{\Theta}\;\mathbb{E}\Biggl[\Bigl(\Phi_{\text{ICNN}}\Bigl(x,\Psi^{2}\bigl(\tfrac{1}{|S|}\sum_{s=1}^{|S|}\Psi^{1}(\xi_{s})\bigr)\Bigr)-\mathbb{E}_{\xi}[Q(x,\xi)]\Bigr)^{2}\Biggr],$ (17)
where $\Theta=\{W_{k},S_{k},b_{k}\}\cup\theta_{\Psi^{1}}\cup\theta_{\Psi^{2}}$ represents the complete set of trainable parameters of $\Phi_{\text{ICNN}}$, $\Psi^{1}$, and $\Psi^{2}$. Importantly, once trained, the model generalises to arbitrary second-stage scenario sets over the same problem structure without requiring retraining, eliminating scenario set-specific training overhead.
@@ 4.3 Surrogate embedding
The ICNN architecture enables an exact convex optimisation formulation to be used instead of Neur2SP’s MIP formulation. Building on the exact inference equivalence (Section 2.3), we reformulate the 2SP as:
$\underset{\displaystyle x,z_{k}}{\mathrm{min}}\quad c^{\intercal}x+z_{K}$ (18a)
@@
$z_{k}\geq 0,\quad$ $k=1,\ldots,K$ (18c)
$x\in\mathcal{X}.$ (18d)
At deployment, $\xi_{\lambda}$ is precomputed offline from a given scenario set, ensuring the problem remains convex in $x$.
@@ 5.1 Experimental setup
All experiments were conducted on the Triton computing cluster maintained by the Aalto Science-IT project, utilising an Intel Xeon Gold 6148 CPU and an NVIDIA Tesla H100 GPU with 64 GB of RAM. Our code extends Neur2SP’s codebase [khalil_neur2sp_2024] by augmenting it to include all necessary ICNN-related functionalities. The implementation uses Python 3.11.5 with Gurobi 11.0.2 [gurobi_optimization_llc_gurobi_2024] as the optimisation solver, PyTorch 2.0.1 [paszke_pytorch_2019] and scikit-learn 1.3.0 [pedregosa_scikit_2011] for NN training. Full reproducibility materials are provided at https://github.com/Lycle/ICNN2SP.
We evaluate our method on three canonical 2SP problems which are also taken as benchmark problems in Neur2SP: the capacitated facility location problem (CFLP) [cornuejols_comparison_1991], the stochastic server location problem (SSLP) [ntaimo_million-variable_2005], and the investment problem (INVP) [schultz_solving_1998]. Problem instances are denoted by type followed by structural parameters, with the final parameter indicating scenario count. Detailed descriptions are provided in C. The EF serves as the ground-truth baseline (considering a 3-hour time limit). We compare both ICNN-enhanced 2SP and Neur2SP against EF to contextualise performance.
To ensure direct comparability with Neur2SP, we replicate their instance generation protocol, including uncertainty distributions and scenario sampling methods. We generate 5,000 first-stage solutions by uniformly sampling the feasible region. For each solution, the expected recourse cost is approximated by solving its second-stage problem for a random subset of scenarios, with the number of scenarios per solution randomly selected up to 100, which balances computational efficiency with the diversity of scenarios. Further details on data generation are provided in D. For both methods, we conduct a random search over 100 hyperparameter configurations for NN architecture selection, detailed in E. To mitigate solver-induced variability, Gurobi’s parameters are fixed across all runs, as detailed in F.
@@ 5.2 Results and discussion
Tables 1 to 7 report our experimental results. We first examine training behaviour in terms of training time and validation accuracy (Table 1), then compare solving performance against EF using gap statistics and solving times (Tables 24). To complement these relative performance indicators, Tables 57 list the underlying objective values for all methods. In addition, we include results for the SSLP instances from the Stochastic Integer Programming Test Problem Library (SIPLIB) [ahmed_siplib_2015] in G. We also analyse the variable counts in H to compare formulation size.
5.2.1 Training performance
Both methods share the same structural design for Scenario Encoding ($\Psi^{1}$ and $\Psi^{2}$) while differing in the architectural choice of Decision Mapping ($\Phi$): our method uses ICNNs while Neur2SP employs standard ReLU NNs, which constitutes the primary source of training performance differences. As shown in Table 1, both methods exhibit training times within the same order of magnitude, with ICNNs requiring only marginally longer training due to their convexity constraints. This modest overhead is counterbalanced by their comparable or superior validation accuracy, achieving lower mean absolute error (MAE) in 7 out of 10 problem instances. These empirical findings are consistent with the convexity properties analysed in Section 3, as the observed error patterns suggest that all benchmark problems exhibit convex or near-convex recourse behaviour, allowing ICNNs to capture the underlying structure effectively. These results demonstrate that the convexity constraints do not significantly compromise the expressive power of the NN while presenting mathematical properties that enable more efficient downstream optimisation.
@@ 5.2.1 Training performance
5.2.2 Optimisation results
[…]
Table 2: CFLP results (averaged over 11 instances per row; EF averages include only instances where EF found feasible solutions within 3h). Gap: relative objective difference (%) of ICNN-enhanced 2SP (ICNN-EF) and Neur2SP (NN-EF) compared to EF; negative values indicate better solutions for minimisation. Solving time: runtime (s); $\sim$10,800 indicates time limit reached. EF time to match (s): average time for EF over instances when it achieved solution quality matching ICNN-enhanced 2SP (ICNN) and Neur2SP (NN); “-” indicates that EF did not match within the time limit. Bold indicates better performance.
Gap (%) Solving time (s) EF time to match (s)
Problem ICNN-EF NN-EF ICNN NN EF ICNN NN
CFLP_10_10_100 0.57 0.51 0.29 0.39 8,491.85 3,801.86 5,281.23
CFLP_10_10_500 0.15 0.00 0.29 0.39 10,800.05 6,363.94 7,456.05
CFLP_10_10_1000 -0.02 -0.27 0.29 0.39 10,800.29 5,946.09 6,331.67
CFLP_25_25_100 2.19 3.37 0.39 10.31 10,800.05 327.97 235.64
CFLP_25_25_500 1.74 2.63 0.40 10.37 10,800.12 1,357.47 1,291.39
CFLP_25_25_1000 0.51 1.48 0.40 10.07 10,800.33 4,133.03 3,949.87
CFLP_50_50_100 3.36 18.17 0.81 84.07 10,800.63 1,432.75 317.61
CFLP_50_50_500 -9.82 3.06 0.84 84.32 10,800.16 - 5,304.11
CFLP_50_50_1000 -12.82 -0.28 0.84 84.39 10,800.29 - -
Table 3: SSLP results (averaged over 12 instances per row, including one from SIPLIB [ahmed_siplib_2015]). All settings and metric definitions follow Table 2.
Gap(%) Solving time (s) EF time to match (s)
Problem ICNN-EF NN-EF ICNN NN EF ICNN NN
SSLP_5_25_50 0.12 0.12 0.13 0.25 1.58 1.36 1.36
SSLP_5_25_100 0.02 0.02 0.12 0.25 5.18 5.01 5.01
SSLP_15_45_5 2.65 18.09 0.16 0.20 1.94 1.15 0.11
SSLP_15_45_10 2.57 18.38 0.18 0.20 32.74 8.51 0.20
SSLP_15_45_15 2.11 18.72 0.20 0.21 605.22 538.02 0.40
SSLP_10_50_50 0.00 0.00 0.27 0.76 9,007.88 12.28 12.28
SSLP_10_50_100 0.00 0.00 0.26 0.76 10,800.04 41.49 41.49
SSLP_10_50_500 0.00 0.00 0.28 0.77 10,800.21 1,570.14 1,570.14
SSLP_10_50_1000 -2.22 -2.22 0.28 0.77 10,803.79 - -
SSLP_10_50_2000 -102.68 -102.68 0.29 0.78 10,800.14 - -
@@ 5.2.2 Optimisation results
Gap (%) Solving time (s) EF time to match (s)
Problem ICNN-EF NN-EF ICNN NN EF ICNN NN
INVP_B_E_4 11.06 10.80 0.06 0.15 0.00 0.00 0.00
INVP_B_E_9 5.60 4.31 0.07 0.10 0.01 0.00 0.00
INVP_B_E_36 3.28 4.58 0.06 0.10 0.05 0.01 0.01
INVP_B_E_121 3.07 2.85 0.06 0.09 1.32 0.01 0.01
INVP_B_E_441 3.10 3.18 0.06 0.10 125.60 0.05 0.05
INVP_B_E_1681 0.98 0.21 0.07 0.10 10,800.00 0.36 4,791.19
INVP_B_E_10000 -1.59 -1.64 0.07 0.11 10,808.33 - -
INVP_B_H_4 8.81 8.81 0.06 0.08 0.01 0.00 0.00
INVP_B_H_9 5.04 5.04 0.07 0.09 0.01 0.01 0.01
INVP_B_H_36 1.61 1.61 0.07 0.08 1.18 0.09 0.09
INVP_B_H_121 1.77 1.77 0.06 0.08 60.09 0.06 0.06
INVP_B_H_441 2.13 2.13 0.06 0.09 322.05 2.60 2.60
INVP_B_H_1681 0.69 0.69 0.06 0.09 10,800.01 886.69 886.69
INVP_B_H_10000 -2.72 -2.72 0.07 0.10 10,808.33 - -
INVP_I_E_4 4.78 17.32 0.05 0.08 0.00 0.00 0.00
INVP_I_E_9 1.07 7.01 0.04 0.07 0.01 0.01 0.00
INVP_I_E_36 0.37 3.98 0.06 0.07 0.03 0.03 0.01
INVP_I_E_121 0.17 2.88 0.06 0.08 0.39 0.20 0.01
INVP_I_E_441 0.11 3.22 0.04 0.07 39.73 1.06 0.05
INVP_I_E_1681 0.42 3.39 0.05 0.07 10,800.00 4.18 0.64
INVP_I_E_10000 0.22 3.50 0.05 0.08 10,800.09 670.21 465.22
INVP_I_H_4 13.78 13.78 0.07 0.06 0.01 0.00 0.00
INVP_I_H_9 5.85 9.12 0.05 0.07 0.01 0.01 0.01
INVP_I_H_36 3.69 4.97 0.06 0.07 0.40 0.02 0.02
INVP_I_H_121 3.75 4.01 0.05 0.07 7.27 0.06 0.06
INVP_I_H_441 1.27 3.15 0.05 0.07 10,800.00 761.23 0.30
INVP_I_H_1681 -0.41 -0.34 0.05 0.07 10,800.01 - -
INVP_I_H_10000 0.27 0.25 0.06 0.08 10,800.03 1,151.03 1,151.03
Tables 57 provide the corresponding objective values behind the gap statistics reported above. For each problem class, the tables list the EF baseline objectives and the true and approximate objectives induced by the surrogate-based formulations. These results confirm that the ICNN-enhanced approach consistently achieves solution quality on par with conventional NN approaches, while retaining the computational advantages highlighted in Tables 24.